# Exploring Weight Balancing on Long-Tailed Recognition Problem

### v5

@@ Abstract
Recognition problems in long-tailed data, in which the sample size per class is heavily skewed, have gained importance because the distribution of the sample size per class in a dataset is generally exponential unless the sample size is intentionally adjusted. Various methods have been devised to address these problems. Recently, weight balancing, which combines well-known classical regularization techniques with two-stage training, has been proposed. Despite its simplicity, it is known for its high performance compared with existing methods devised in various ways. However, there is a lack of understanding as to why this method is effective for long-tailed data. In this study, we analyze weight balancing by focusing on neural collapse and the cone effect at each training stage and found that it can be decomposed into an increase in Fisher’s discriminant ratio of the feature extractor caused by weight decay and cross entropy loss and implicit logit adjustment caused by weight decay and class-balanced loss. Our analysis enables the training method to be further simplified by reducing the number of training stages to one while increasing accuracy.
@@ 1 Introduction
Datasets with an equal number of samples per class, such as MNIST (Lecun et al., 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), are often used, when we evaluate classification models and training methods in machine learning. However, it is empirically known that the size distribution in the real world often shows a type of exponential distribution called Pareto distribution (Reed, 2001), and the same is true for the number of per-class samples in classification problems (Li et al., 2017; Spain and Perona, 2007). Such distributions are called long-tailed data due to the shape of the distribution since some classes (head classes) are often sampled and many others (tail classes) are not sampled very often. Long-tailed recognition (LTR) is used to attempt to improve the accuracy of classification models on uniform distribution when training data shows such a distribution. There is a problem in LTR that the head classes have large sample size; thus, the output is biased toward them. This reduces the overall and tail class accuracy because tail classes make up the majority (Zhang et al., 2021).
Various methods have been developed for LTR, such as class-balanced loss (CB) (Cui et al., 2019), augmenting samples of tail classes (Wang et al., 2021), two-stage learning (Kang et al., 2020), and enhancing feature extractors (Liu et al., 2023; Yang et al., 2022); see Appendix A.2 for more related research. Alshammari et al. (2022) proposed a simple method, called weight balancing (WB), that empirically outperforms previous complex state-of-the-art methods. WB simply combines two classic techniques, weight decay (WD) (Hanson and Pratt, 1989) and MaxNorm (Hinton et al., 2012), with two-stage learning. WD and MaxNorm is known to prevent overlearning (Hinton, 1989); however, it is not known why WB significantly improves the LTR performance.
@@
The above analysis reveals the following useful points: 1. our analysis provides a guideline for the design of learning in LTR; 2. our theorem offers an insight into how to suppress the cone effect (Liang et al., 2022) that negatively affects deep learning classification models; 3. WB can be further simplified by removing the second stage. Specifically, we only need to learn in LTR by WD, feature regularization (FR), and an equiangular tight frame (ETF) classifier for the linear layer to extract features that are more linearly separable and adjusting the norm of classifier’s weights by LA after the training.
@@ 2 Related Work
Papyan et al. (2020) investigated the set of phenomena that occurs when the terminal phase of training, which they termed NC. NC can be briefly described as follows. Feature vectors converge to their class means (NC1); the class means of feature vectors converge to a simplex ETF (Strohmer and Heath, 2003) (NC2); the class means of feature vectors and the corresponding weights of the linear classifier converge in the same direction (NC3); and when making predictions, models converge to predict the class, the mean feature vector of which is closest to the feature in Euclidean distance (NC4).
Papyan et al. (2020) claimed that NC leads to increased generalization accuracy and robustness to adversarial samples. The number of studies have been conducted on the conditions under which NC occurs (Han et al., 2022; Ji et al., 2022; Lu and Steinerberger, 2021; Papyan et al., 2020). Rangamani and Banburski-Fahey (2022) theoretically and empirically showed that the occurrence of NC needs WD.
@@
WD is a regularization method often used for deep neural network models with BN (Ioffe and Szegedy, 2015), which is scale invariant. Zhang et al. (2019) revealed that WD increases the effective learning rate which facilitates regularization in such models. Training of networks containing BN layers with WD is often studied in terms of training dynamics (Li and Arora, 2020; Lobacheva et al., 2021; Wan et al., 2021). Summers and Dinneen (2020) and Kim et al. (2022) studied whether to apply WD to the scaling and shifting parameters of the BN for such networks. WD is also known to have many other positive effects on deep neural networks, e.g., smoothing loss landscape (Li et al., 2018; Lyu et al., 2022), causing NC (Rangamani and Banburski-Fahey, 2022), and making filters sparser (Mehta et al., 2019) and low-ranked (Galanti et al., 2023). WD is also effective for imbalanced data (Alshammari et al., 2022; Thrampoulidis et al., 2022).
Liang et al. (2022) found a phenomenon they termed “cone effect” observed in neural networks. The phenomenon is a tendency in which features from deep neural networks with activation functions are prone to have high cosine similarities. It is widely observed regardless of the modality of data, the structure of models, and whether the models are trained. It also indicates that features from different classes tend to exhibit high cosine similarity. Thus, it has a negative impact on deep neural networks for classification since features with lower inter-class cosine similarity are more linearly separated. We also reconfirmed such phenomena in our experiments (see Sec. 4.2). Our analysis provides guidelines to prevent the effect.
@@ 3 Preliminaries
Multiplicative LA (Kim and Kim, 2020) changes the norm of the per-class weights of the linear layer. The weights of the linear layer corresponding to class $k$, $\mathbf{w}^{\prime}_{k}$ are adjusted to $\frac{1}{{\mathbb{P}(Y=k)}^{\gamma}}\frac{\mathbf{w}_{k}}{\|\mathbf{w}_{k}\|_{2}}$, where $\gamma>0$ is a hyperparameter. Kim and Kim (2020) trained using projected gradient descent so that the weights of the linear layer always satisfy $\forall k,\ \|\mathbf{w}_{k}\|_{2}=1$, but in our study, for comparison, we trained with regular gradient descent, normalized the norm post-hoc, and then adjusted the norm. Additive LA (Menon et al., 2020) is a method to adjust the output to minimize the average per-class error by adding a different constant for each class to the logit at prediction. That is, the logit $\mathbf{z}_{i}$ at prediction is adjusted to $\mathbf{z}^{\prime}_{i}$ as $(\mathbf{z}^{\prime}_{i})_{k}=(\mathbf{z}_{i})_{k}-\tau\log\mathbb{P}(Y=k),$ where $\tau$ is a hyperparameter. LA generally increases the probability of classifying inputs into tail classes.
@@ 4.2 WD and CE degrade inter-class cosine similarities
Table 1 lists the FDRs of the models trained with each method, and Figure 1 shows the cosine similarities of the training features per class. It can be seen that the combination of WD and CE achieves the highest FDR. This fact is consistent with the conclusion that WD is necessary for the NC in balanced datasets, as shown by Rangamani and Banburski-Fahey (2022) and suggests that this may also hold for long-tailed data.
When WD is not applied, the cosine similarity is generally higher even between features of different classes. This is natural because neural networks have the cone effect found in Liang et al. (2022). The cone effect is a phenomenon in which features from a deep neural network tend to have high cosine similarities to each other even if they belong to different classes. However, the methods with WD result in lower cosine similarities between features of different classes and higher cosine similarities between features of the same classes, leading to a higher FDR. This shows that WD prevents the cone effect, as supported by the following theorem.
@@ 5.1 WD and CB perform implicit logit adjustment
where $O$ is a Bachmann-Landau O-notation. For example, the notation in the form $O\left(\frac{1}{x}\right)$ indicates that the term is at most a constant multiple of $\frac{1}{x}$ if $x$ is sufficiently large. Note that $\frac{\overline{N}}{\lambda N}$ is independent of $k$. This theorem states that if the number of classes or the imbalance factor is sufficiently large, and if NC has occurred in the first stage, there exists linear layer weights that are constant multiples of the corresponding features and sufficiently close to the stationary point in the optimization. In other words, if the previous conditions are satisfied, the norm of the linear layer is proportional to the norm of the corresponding per-class mean feature. We show that the norm of training features are larger for Few classes in Sec. 4.5. Thus, the linear layer is trained to be larger for the Few classes. Besides, the features and the corresponding linear layers’ weights are aligned when NC occurs, so the orientation of the linear layers’ weights hardly changes during this training. This is the same operation as when multiplicative LA makes adjustments to long-tailed data, increasing the norm of the linear layer for the tail class (Kim and Kim, 2020). Indeed, if there exists some $c_{0}$ and $\gamma_{0}$, and for all $k$, $\left\|\bm{\mu}_{k}\right\|_{2}=c_{0}\mathbb{P}(Y=k)^{\gamma_{0}}$ holds, then the second-stage training is equivalent to multiplicative LA. In this theorem, we do not consider MaxNorm because we found that MaxNorm makes the norm of the weights in the linear layer closer to the same value before training, only facilitating convergence. For more details, see Appendix A.5.
@@ 5.2 Is two-stage learning really necessary?
As for the bottom three methods, we also compared with additive LA (Menon et al., 2020), with multiplicative LA (Kim and Kim, 2020), and without LA. To ensure that the model correctly classifies both samples of the head classes and tail classes, we consider the average of the per-class accuracy of all classes and each group, i.e., Many, Medium, and Few.
@@ A.1 Acronym and Notation
BN batch normalization [Ioffe and Szegedy, 2015]
CB class-balanced loss [Cui et al., 2019]
@@
FDR Fisher’s discriminant ratio [Fisher, 1936]
@@
LA logit adjustment [Kim and Kim, 2020, Menon et al., 2020]
@@
NC neural collapse [Papyan et al., 2020]
ResBlock residual block [He et al., 2016]
@@
WB weight balancing [Alshammari et al., 2022]
@@ A.2.1 Long-tailed recognition
There are three main approaches to LTR; “Class Re-balancing, Information Augmentation, and Module Improvement” [Zhang et al., 2021]. “Class Re-balancing” adjusts the imbalance in the number of samples per class at various stages to prevent accuracy deterioration. It includes logit adjustment (LA) [Kim and Kim, 2020, Menon et al., 2020] and balancing the loss function such as CB and dynamic semantic-scale-balanced loss [Ma et al., 2022]. “Information Augmentation” prevents the accuracy from degrading by supplementing the information of the tail classes that lacks the number of samples [Chu et al., 2020, Liu et al., 2020, Wang et al., 2021]. “Module Improvement” improves accuracy by increasing the performance of each module of the network individually, e.g., training feature extractors and classifiers separately [Kang et al., 2020], fixing the linear layer to the ETF classifier for the better feature extractor [Yang et al., 2022], and regularizing to occur NC [Liu et al., 2023]. In recent years, as Kang et al. [2023] found contrastive learning [Hadsell et al., 2006] is effective for imbalanced data, many methods have been proposed to use it [Kang et al., 2023, Li et al., 2022, Ma et al., 2021, Tian et al., 2022]. Ma et al. [2021], Long et al. [2022], and Tian et al. [2022] also include vision-language models such as CLIP [Radford et al., 2021] and leverage text features to improve accuracy. However, these usually have problems such as slow convergence and complex models [Liu et al., 2023]. WB is a combination of “Class Re-balancing” and “Module Improvement”. Table 8 compares our simplification of WB with typical existing methods, including those that have achieved SOTA. While many methods use complex innovations, we aimed to improve on a simple structure.
@@
Two-stage learning [Kang et al., 2020] is a method for improving accuracy by dividing LTR training into two stages: feature extractor training and classifier training. It is used in numerous methods [Cao et al., 2019, Ma et al., 2021, 2022, Li et al., 2022, Tian et al., 2022, Liu et al., 2023, Kang et al., 2023], including WB, because of its simple but significant improvement in accuracy. Note that since our work analyzes WB, the formulation of two-stage learning is based on WB. For example, the classifier weights are initialized randomly at the start of the second stage of training in Kang et al. [2020] but not in WB.
Table 8: Comparison of existing LTR methods with our simplified method. A symbol  indicates that the component is required, - indicates that the component is not required. We regard the deferred re-balancing optimization schedule [Cao et al., 2019] as two-stage training. Our method does not require complex technique like many other methods.
@@
Kim and Kim [2020]
Menon et al. [2020]
Yang et al. [2022]
Ma et al. [2022]
Liu et al. [2023]
Alshammari et al. [2022]
Cao et al. [2019]
Kang et al. [2023]
Li et al. [2022]
Long et al. [2022]
Ma et al. [2021]
Tian et al. [2022]
@@
Number of training stages 1 1 1 3 2 2 2 2 2 $*{}^{*}$ 2 $*{}^{*}$ 1
@@
Training of linear layer - - [^2] This method does not require linear layers.  [^3] This method requires the LGR head to be trained [Tian et al., 2022]. -

### v7

@@ Abstract
Recognition problems in long-tailed data, in which the sample size per class is heavily skewed, have gained importance because the distribution of the sample size per class in a dataset is generally exponential unless the sample size is intentionally adjusted. Various methods have been devised to address these problems. Recently, weight balancing, which combines well-known classical regularization techniques with two-stage training, has been proposed. Despite its simplicity, it is known for its high performance compared with existing methods devised in various ways. However, there is a lack of understanding as to why this method is effective for long-tailed data. In this study, we analyze weight balancing by focusing on neural collapse and the cone effect at each training stage and found that it can be decomposed into an increase in Fisher’s discriminant ratio of the feature extractor caused by weight decay and cross entropy loss and implicit logit adjustment caused by weight decay and class-balanced loss. Our analysis enables the training method to be further simplified by reducing the number of training stages to one while increasing accuracy. Code is available at https://github.com/HN410/Exploring-Weight-Balancing-on-Long-Tailed-Recognition-Problem.
@@ 1 Introduction
Datasets with an equal number of samples per class, such as MNIST (Lecun et al., 1998) and CIFAR100 (Krizhevsky, 2009), are often used, when we evaluate classification models and training methods in machine learning. However, it is empirically known that the size distribution in the real world often shows a type of exponential distribution called Pareto distribution (Reed, 2001), and the same is true for the number of per-class samples in classification problems (Li et al., 2017; Spain & Perona, 2007). Such distributions are called long-tailed data due to the shape of the distribution since some classes (head classes) are often sampled and many others (tail classes) are not sampled very often. Long-tailed recognition (LTR) is used to attempt to improve the accuracy of classification models on uniform distribution when training data shows such a distribution. There is a problem in LTR that the head classes have large sample size; thus, the output is biased toward them. This reduces the overall and tail class accuracy because tail classes make up the majority (Zhang et al., 2021).
Various methods have been developed for LTR, such as class-balanced loss (CB) (Cui et al., 2019), augmenting samples of tail classes (Wang et al., 2021), two-stage learning (Kang et al., 2020), and enhancing feature extractors (Liu et al., 2023; Yang et al., 2022); see Appendix A.2 for more related research. Alshammari et al. (2022) proposed a simple method, called weight balancing (WB), that empirically outperforms previous complex state-of-the-art methods. WB simply combines two classic techniques, weight decay (WD) (Hanson & Pratt, 1989) and MaxNorm (Hinton et al., 2012), with two-stage learning. WD and MaxNorm is known to prevent overlearning (Hinton, 1989); however, it is not known why WB significantly improves the LTR performance.
@@
The above analysis is useful in the following points: 1. it provides a guideline for the design of learning in LTR; 2. our theorem offers an insight into how to suppress the cone effect (Liang et al., 2022) that negatively affects deep learning classification models; 3. WB can be further simplified by removing the second stage. Specifically, we only need to learn in LTR by WD, feature regularization (FR), and an equiangular tight frame (ETF) classifier for the linear layer to extract features that are more linearly separable and adjusting the norm of classifier’s weights by LA after the training.
@@ 2 Related Work
Papyan et al. (2020) investigated the set of phenomena that occurs when the terminal phase of training, which they termed NC. NC can be briefly described as follows. Feature vectors converge to their class means (NC1); the class means of feature vectors converge to a simplex ETF (Strohmer & Heath, 2003) (NC2); the class means of feature vectors and the corresponding weights of the linear classifier converge in the same direction (NC3); and when making predictions, models converge to predict the class, the mean feature vector of which is closest to the feature in Euclidean distance (NC4).
Papyan et al. (2020) claimed that NC leads to increased generalization accuracy and robustness to adversarial samples. The number of studies have been conducted on the conditions under which NC occurs (Han et al., 2022; Ji et al., 2022; Lu & Steinerberger, 2021; Papyan et al., 2020). Rangamani & Banburski-Fahey (2022) theoretically and empirically showed that the occurrence of NC needs WD.
@@
WD is a regularization method often used for deep neural networks (DNNs) with BN (Ioffe & Szegedy, 2015), which is scale invariant. Zhang et al. (2019) revealed that WD increases the effective learning rate which facilitates regularization in such models. Training of networks containing BN layers with WD is often studied in terms of training dynamics (Li & Arora, 2020; Lobacheva et al., 2021; Wan et al., 2021). Summers & Dinneen (2020) and Kim et al. (2022) studied whether to apply WD to the scaling and shifting parameters of the BN for such networks. WD is also known to have many other positive effects on DNNs, e.g., smoothing loss landscape (Li et al., 2018; Lyu et al., 2022), causing NC (Rangamani & Banburski-Fahey, 2022), and making filters sparser (Mehta et al., 2019) and low-ranked (Galanti et al., 2023). WD is also effective for imbalanced data (Alshammari et al., 2022; Thrampoulidis et al., 2022).
Liang et al. (2022) found a phenomenon they termed “cone effect” observed in neural networks. The phenomenon is a tendency in which features from DNNs with activation functions are prone to have high cosine similarities. It is widely observed regardless of the modality of data, the structure of models, and whether the models are trained. It also indicates that features from different classes tend to exhibit high cosine similarity. Thus, it has a negative impact on DNNs for classification since features with lower inter-class cosine similarity are more linearly separated. We also reconfirmed such phenomena in Sec. 4.2. Our analysis provides guidelines to prevent the effect.
@@ 3 Preliminaries
Multiplicative LA (Kim & Kim, 2020) changes the norm of the per-class weights of the linear layer. The weights of the linear layer corresponding to class $k$, $\mathbf{w}^{\prime}_{k}$ are adjusted to $\frac{1}{{\mathbb{P}(Y=k)}^{\gamma}}\frac{\mathbf{w}_{k}}{\|\mathbf{w}_{k}\|_{2}}$, where $\gamma>0$ is a hyperparameter. Kim & Kim (2020) trained using projected gradient descent so that the weights of the linear layer always satisfy $\forall k,\ \|\mathbf{w}_{k}\|_{2}=1$, but in our study, for comparison, we trained with regular gradient descent, normalized the norm post-hoc, and then adjusted the norm. Additive LA (Menon et al., 2020) is a method to adjust the output to minimize the average per-class error by adding a different constant for each class to the logit at prediction. That is, the logit $\mathbf{z}_{i}$ at prediction is adjusted to $\mathbf{z}^{\prime}_{i}$ as $(\mathbf{z}^{\prime}_{i})_{k}=(\mathbf{z}_{i})_{k}-\tau\log\mathbb{P}(Y=k),$ where $\tau$ is a hyperparameter. LA generally increases the probability of classifying inputs into tail classes.
@@ 4.2 WD and CE degrade inter-class cosine similarities
Table 1 lists the FDRs of the models trained with each method, and Figure 1 shows the cosine similarities of the training features per class. It can be seen that the combination of WD and CE achieves the highest FDR. This fact is consistent with the conclusion that WD is necessary for the NC in balanced datasets, as shown by Rangamani & Banburski-Fahey (2022) and suggests that this may also hold for long-tailed data.
When WD is not applied, the cosine similarity is generally higher even between features of different classes. This is natural because neural networks have the cone effect found in Liang et al. (2022). The cone effect is a phenomenon in which features from a DNN tend to have high cosine similarities to each other even if they belong to different classes. However, the methods with WD result in lower cosine similarities between features of different classes and higher cosine similarities between features of the same classes, leading to a higher FDR. This shows that WD prevents the cone effect, as supported by the following theorem.
@@ 5.1 WD and CB perform implicit logit adjustment
where $O$ is a Bachmann-Landau O-notation. For example, the notation in the form $O\left(\frac{1}{x}\right)$ indicates that the term is at most a constant multiple of $\frac{1}{x}$ if $x$ is sufficiently large. Note that $\frac{\overline{N}}{\lambda N}$ is independent of $k$. This theorem states that if the number of classes or the imbalance factor is sufficiently large, and if NC has occurred in the first stage, there exists linear layer weights that are constant multiples of the corresponding features and sufficiently close to the stationary point in the optimization. In other words, if the previous conditions are satisfied, the norm of the linear layer is proportional to the norm of the corresponding per-class mean feature. We show that the norm of training features are larger for Few classes in Sec. 4.5. Thus, the linear layer is trained to be larger for the Few classes. Besides, the features and the corresponding linear layers’ weights are aligned when NC occurs, so the orientation of the linear layers’ weights hardly changes during this training. This is the same operation as when multiplicative LA makes adjustments to long-tailed data, increasing the norm of the linear layer for the tail class (Kim & Kim, 2020). Indeed, if there exists some $c_{0}$ and $\gamma_{0}$, and for all $k$, $\left\|\bm{\mu}_{k}\right\|_{2}=c_{0}\mathbb{P}(Y=k)^{\gamma_{0}}$ holds, then the second-stage training is equivalent to multiplicative LA. In this theorem, we do not consider MaxNorm because we found that MaxNorm makes the norm of the weights in the linear layer closer to the same value before training, only facilitating convergence. For more details, see Appendix A.5.
@@ 5.2 Is two-stage learning really necessary?
As for the bottom three methods, we also compared with additive LA (Menon et al., 2020), with multiplicative LA (Kim & Kim, 2020), and without LA. To ensure that the model correctly classifies both samples of the head classes and tail classes, we consider the average of the per-class accuracy of all classes and each group, i.e., Many, Medium, and Few.
@@ 6 Conclusion
Acknowledgements
We would like to express my sincere gratitude to the members of Issei Sato Laboratory for their invaluable discussions. We are also grateful to the reviewers for their diligent critique, which has significantly improved the quality of this paper. This work was supported by JSPS KAKENHI Grant Number 20H04239 Japan.
@@ A.1 Acronym and Notation
BN batch normalization (Ioffe & Szegedy, 2015)
CB class-balanced loss (Cui et al., 2019)
@@
FDR Fisher’s discriminant ratio (Fisher, 1936)
@@
LA logit adjustment (Kim & Kim, 2020; Menon et al., 2020)
@@
NC neural collapse (Papyan et al., 2020)
ResBlock residual block (He et al., 2016)
@@
WB weight balancing (Alshammari et al., 2022)
@@ A.2.1 Long-tailed recognition
There are three main approaches to LTR; “Class Re-balancing, Information Augmentation, and Module Improvement” (Zhang et al., 2021). “Class Re-balancing” adjusts the imbalance in the number of samples per class at various stages to prevent accuracy deterioration. It includes logit adjustment (LA) (Kim & Kim, 2020; Menon et al., 2020) and balancing the loss function such as CB and dynamic semantic-scale-balanced loss (Ma et al., 2022). “Information Augmentation” prevents the accuracy from degrading by supplementing the information of the tail classes that lacks the number of samples (Chu et al., 2020; Liu et al., 2020; Wang et al., 2021). “Module Improvement” improves accuracy by increasing the performance of each module of the network individually, e.g., training feature extractors and classifiers separately (Kang et al., 2020), fixing the linear layer to the ETF classifier for the better feature extractor (Yang et al., 2022), and regularizing to occur NC (Liu et al., 2023). In recent years, as Kang et al. (2023) found contrastive learning (Hadsell et al., 2006) is effective for imbalanced data, many methods have been proposed to use it (Kang et al., 2023; Li et al., 2022; Ma et al., 2021; Tian et al., 2022). Ma et al. (2021), Long et al. (2022), and Tian et al. (2022) also include vision-language models such as CLIP (Radford et al., 2021) and leverage text features to improve accuracy. However, these usually have problems such as slow convergence and complex models (Liu et al., 2023). WB is a combination of “Class Re-balancing” and “Module Improvement”. Table 8 compares our simplification of WB with typical existing methods, including those that have achieved SOTA. While many methods use complex innovations, we aimed to improve on a simple structure.
@@
Two-stage learning (Kang et al., 2020) is a method for improving accuracy by dividing LTR training into two stages: feature extractor training and classifier training. It is used in numerous methods (Cao et al., 2019; Ma et al., 2021; 2022; Li et al., 2022; Tian et al., 2022; Liu et al., 2023; Kang et al., 2023), including WB, because of its simple but significant improvement in accuracy. Note that since our work analyzes WB, the formulation of two-stage learning is based on WB. For example, the classifier weights are initialized randomly at the start of the second stage of training in Kang et al. (2020) but not in WB.
Table 8: Comparison of existing LTR methods with our simplified method. A symbol  indicates that the component is required, - indicates that the component is not required. We regard the deferred re-balancing optimization schedule (Cao et al., 2019) as two-stage training. Our method does not require complex technique like many other methods.
@@
Kim & Kim (2020)
Menon et al. (2020)
Yang et al. (2022)
Ma et al. (2022)
Liu et al. (2023)
Alshammari et al. (2022)
Cao et al. (2019)
Kang et al. (2023)
Li et al. (2022)
Long et al. (2022)
Ma et al. (2021)
Tian et al. (2022)
@@
Number of training stages 1 1 1 3 2 2 2 2 2 1 2 2 1
@@
Training of linear layer - - [^2] This method does not require linear layers.  [^3] This method requires the LGR head to be trained (Tian et al., 2022). -